OpenStudy (anonymous):

The perimeter of a rectangle is 56 inches. The length is 4 more than 3 times the width. What are the dimensions of the rectangle? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

x = 3y+4 x = lenght y = width 2(3y+4) + 2y = 56 8y = 48 y = 6 x = 22

OpenStudy (anonymous):

i dont understand here are my options L+W=56 L=3W+4 L+W=56 W+4=3L 2L+2W=56 W=3L+4 2L+2W=56 L=3W+4

OpenStudy (anonymous):

A (L+W=56 L=3W+4 ) B (L+W=56 W+4=3L) C (2L+2W=56 W=3L+4) D (2L+2W=56 L=3W+4)

OpenStudy (anonymous):

D.